Please show formula to find the answer1) You have just won the lottery. At the end of each of the next 20 years, you will receive a payment of $50,000. If the cost3) I now have $250,000 in the bank. At the end of each of the next 20 years, I withdraw $15,000 to live on. If I earn 8 perce
Answer #1

1.
=50000/10%*(1-1/1.1^20)=425678.185987928

2.
=14000/10%*1.1=154000

3.
=(250000-15000/8%*(1-1/1.08^20))*1.08^20=478809.821490582

4.
=15000000/6%*(1-1/1.06^7)=83735721.594416